OpenStudy (itz_sid):

A spring has a natural length of 10 cm. If a 26-N force is required to keep it stretched to a length of 22 cm, how much work W is required to stretch it from 10 cm to 16 cm? (Round your answer to two decimal places.)

OpenStudy (itz_sid):

Here is my work. But i got the problem wrong :/

OpenStudy (518nad):

work = integral f.dr

OpenStudy (518nad):

first of all did u convert everything to meters

OpenStudy (518nad):

26 = -k((22-10)/100) k=2600/12 =650/3

OpenStudy (518nad):

hello?`

OpenStudy (518nad):

w=kx^2/2 = 650/3 (0.06)^2/2 =650/6 * 6^2/100^2 = 6.5*6/100 = 0.39 Joules

OpenStudy (itz_sid):

Yea i did convert. :/

OpenStudy (itz_sid):

Where did you get -k from?

OpenStudy (itz_sid):

Oh I get it... Thanks!